# Inexact Moreau Envelope Lagrangian Method for Non-Convex Constrained Optimization under Local Error Bound Conditions on Constraint Functions

### v1

@@ 3 Preliminaries
$\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}},\boldsymbol{\lambda}):=f({\mathbf{x}})+\sum_{i=1}^{m}\lambda_{i}g_{i}({\mathbf{x}})+\frac{p}{2}\|{\mathbf{x}}-{\mathbf{z}}\|^{2}.$ (16)
Similar to the notations in [50, 51, 52], we define
$d(\boldsymbol{\lambda},{\mathbf{z}}):=\min_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}},\boldsymbol{\lambda}),\quad{\mathbf{x}}(\boldsymbol{\lambda},{\mathbf{z}}):=\mathop{\mathrm{arg\,min}{}}_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}},\boldsymbol{\lambda}).$ (17)
By the strong duality (see, e.g. [40, Section 28-30]), it holds that
@@ 4 Inexact Moreau Envelope Lagrangian Method
We present the inexact Moreau envelope Lagrangian (iMELa) method in Algorithm 1, for finding an $\epsilon$-KKT point of (1). At iteration $t$, we first obtain $\boldsymbol{\lambda}^{(t+1)}$ by performing a projected gradient ascent step with respect to $\boldsymbol{\lambda}$ on function $\mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda})$ at $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{(t)}$ with a step-size $\tau_{t}$. Then, given the convex subproblem $\min_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$, the optimal solution of which denoted by $\tilde{{\mathbf{x}}}^{(t+1)}:={\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t)})$ in the rest of the paper for simplicity of notation, we approximately solve it by computing its $\epsilon_{t}$-stationary point with $\epsilon_{t}\in(0,1)$, which is a point ${\mathbf{x}}^{(t+1)}\in\mathcal{X}$ satisfying
@@
The point ${\mathbf{x}}^{(t+1)}$ can be found by various FOMs, for example, the algorithms in [35, Section 5.1.1] and [36, (2.2.22)]. Moreover, condition (19) can be numerically verified and thus used as a stopping condition for solving the subproblem, provided that $\mathcal{N}_{\mathcal{X}}({\mathbf{x}})$ has a simple structure. At last, we update ${\mathbf{z}}$ by taking a convex combination of ${\mathbf{z}}^{(t)}$ and ${\mathbf{x}}^{(t+1)}$ with a weight $\theta_{t}\in[0,1]$, which ensures ${\mathbf{z}}^{(t)}\in\mathcal{X},\,\forall\,t\geq 0$.
@@
3:  for iteration $t=0,1,\cdots,T-1$ do
4:     Set $\lambda_{i}^{(t+1)}=[\lambda_{i}^{(t)}+\tau_{t}\cdot g_{i}({\mathbf{x}}^{(t)})]_{+},\;\forall\,i\in[m]$.
5:     Compute ${\mathbf{x}}^{(t+1)}\approx\tilde{{\mathbf{x}}}^{(t+1)}:=\mathop{\mathrm{arg\,min}{}}_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$ that satisfies (19).
6:     Set ${\mathbf{z}}^{(t+1)}={\mathbf{z}}^{(t)}+\theta_{t}\cdot({\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)})$.
@@
Condition (19) holds if and only if there exists ${\mathbf{v}}^{(t+1)}\in\mathbb{R}^{n}$ such that $\|{\mathbf{v}}^{(t+1)}\|\leq\epsilon_{t}$ and $\boldsymbol{\xi}^{(t+1)}\in\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(t+1)})$ such that
This implies that for any ${\mathbf{x}}\in\mathcal{X}$,
In addition, since $\mathcal{L}_{p}$ is $(p-L)$-strongly convex in ${\mathbf{x}}$, we have
$\leavevmode\nobreak\ \mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
$\leq$ $-\langle\nabla_{\mathbf{x}}\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)}),\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}^{(t+1)}\rangle-\frac{p-L}{2}\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}^{(t+1)}\|^{2}$
$\leq$ $\leavevmode\nobreak\ \langle{\mathbf{v}}^{(t+1)},\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}^{(t+1)}\rangle-\frac{p-L}{2}\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}^{(t+1)}\|^{2}$
$\leq$ $\leavevmode\nobreak\ \frac{1}{2(p-L)}\|{\mathbf{v}}^{(t+1)}\|^{2}\leq\frac{\epsilon_{t}^{2}}{2(p-L)},$ (22)
where the second inequality is by (21) at ${\mathbf{x}}=\tilde{{\mathbf{x}}}^{(t+1)}$ and the third one is by Young’s inequality.
The convergence analysis for Algorithm 1 relies on a uniform upper bound on $\{\|\boldsymbol{\lambda}^{(t)}\|\}_{t\geq 0}$, which we present below.
@@
The sequence $\{\boldsymbol{\lambda}^{(t)}\}_{t\geq 0}$ generated by Algorithm 1 satisfies
$\|\boldsymbol{\lambda}^{(t)}\|\leq M_{\boldsymbol{\lambda}}:=\max\Big{\{}2\overline{\tau}B_{g},\frac{2\overline{\tau}(C_{\boldsymbol{\lambda}}+\overline{\tau}B_{g}^{2})}{2\underline{\tau}\min_{i\in[m]}[-g_{i}({\mathbf{x}}_{\textup{feas}})]}\Big{\}},\leavevmode\nobreak\ \forall\,t\geq 0,$ (23)
@@
Proof.
$\langle\boldsymbol{\lambda}^{(t+1)},{\mathbf{g}}({\mathbf{x}}^{(t+1)})-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\leq$ $\leavevmode\nobreak\ \textstyle\sum_{i=1}^{m}\langle\lambda_{i}^{(t+1)}\nabla g_{i}({\mathbf{x}}^{(t+1)}),{\mathbf{x}}^{(t+1)}-{\mathbf{x}}_{\text{feas}}\rangle$
$=$ $\leavevmode\nobreak\ \langle\nabla_{\mathbf{x}}\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})+{\mathbf{v}}^{(t+1)},{\mathbf{x}}^{(t+1)}-{\mathbf{x}}_{\text{feas}}\rangle$
$\leq$ $\leavevmode\nobreak\ \langle\nabla f({\mathbf{x}}^{(t+1)})+p({\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)})+{\mathbf{v}}^{(t+1)},{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}^{(t+1)}\rangle$
$\leq$ $\leavevmode\nobreak\ (B_{f}+pD_{\mathcal{X}}+1)\cdot D_{\mathcal{X}}=C_{\boldsymbol{\lambda}}.$ (25)
where the second inequality is by (24) and the last one is from Assumption 3.1A and D, and the fact that $\|{\mathbf{v}}^{(t+1)}\|\leq\epsilon_{t}<1$. Since $\boldsymbol{\lambda}^{(0)}={\mathbf{0}}$ in Algorithm 1, (25) also holds for $t=-1$. So, $\langle\boldsymbol{\lambda}^{(t)},{\mathbf{g}}({\mathbf{x}}^{(t)})-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\leq C_{\boldsymbol{\lambda}},\leavevmode\nobreak\ \forall\,t\geq 0$, or equivalently,
$\langle\boldsymbol{\lambda}^{(t)},{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle\leq C_{\boldsymbol{\lambda}}-\langle\boldsymbol{\lambda}^{(t)},-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle,\leavevmode\nobreak\ \forall\,t\geq 0.$ (26)
$\|\boldsymbol{\lambda}^{(t)}\|^{2}-2\tau_{t}\langle\boldsymbol{\lambda}^{(t)},-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\leq$ $\leavevmode\nobreak\ \max_{\boldsymbol{\lambda}\geq{\mathbf{0}},\,\|\boldsymbol{\lambda}\|\leq M_{\boldsymbol{\lambda}}}\left\{\|\boldsymbol{\lambda}\|^{2}-2\tau_{t}\langle\boldsymbol{\lambda},-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\right\}$
$=$ $\leavevmode\nobreak\ \max_{\theta\in[0,1],\,\boldsymbol{\lambda}\geq{\mathbf{0}},\,\|\boldsymbol{\lambda}\|=M_{\boldsymbol{\lambda}}}\left\{\|\theta\boldsymbol{\lambda}\|^{2}-2\theta\tau_{t}\langle\boldsymbol{\lambda},-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\right\}.$ (27)
$\|\boldsymbol{\lambda}^{(t)}\|^{2}-2\tau_{t}\langle\boldsymbol{\lambda}^{(t)},-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\leq$ $\leavevmode\nobreak\ \max_{\boldsymbol{\lambda}\geq{\mathbf{0}},\,\|\boldsymbol{\lambda}\|=M_{\boldsymbol{\lambda}}}\left\{\|\boldsymbol{\lambda}\|^{2}-2\tau_{t}\langle\boldsymbol{\lambda},-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\right\}$
$=$ $\leavevmode\nobreak\ M_{\boldsymbol{\lambda}}^{2}-2\tau_{t}M_{\boldsymbol{\lambda}}\cdot\min_{i\in[m]}[-g_{i}({\mathbf{x}}_{\text{feas}})].$ (28)
$\|\boldsymbol{\lambda}^{(t+1)}\|^{2}=$ $\leavevmode\nobreak\ \textstyle\sum_{i=1}^{m}([\lambda_{i}^{(t)}+\tau_{t}\cdot g_{i}({\mathbf{x}}^{(t)})]_{+})^{2}$
$\leq$ $\leavevmode\nobreak\ \textstyle\sum_{i=1}^{m}(\lambda_{i}^{(t)}+\tau_{t}\cdot g_{i}({\mathbf{x}}^{(t)}))^{2}=\|\boldsymbol{\lambda}^{(t)}\|^{2}+2\tau_{t}\langle\boldsymbol{\lambda}^{(t)},{\mathbf{g}}(x^{(t)})\rangle+\tau_{t}^{2}\|{\mathbf{g}}(x^{(t)})\|^{2}$
$\leq$ $\leavevmode\nobreak\ \|\boldsymbol{\lambda}^{(t)}\|^{2}+2\tau_{t}\big{(}C_{\boldsymbol{\lambda}}-\langle\boldsymbol{\lambda}^{(t)},-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\big{)}+\tau_{t}^{2}\|{\mathbf{g}}(x^{(t)})\|^{2}$
$\leq$ $\leavevmode\nobreak\ M_{\boldsymbol{\lambda}}^{2}+2\overline{\tau}C_{\boldsymbol{\lambda}}-2\underline{\tau}M_{\boldsymbol{\lambda}}\cdot\min_{i\in[m]}[-g_{i}({\mathbf{x}}_{\text{feas}})]+\overline{\tau}^{2}B_{g}^{2}\leq M_{\boldsymbol{\lambda}}^{2},$
@@
By (16), Lemma 4.1, and $\|\boldsymbol{\lambda}\|_{1}\leq\sqrt{m}\|\boldsymbol{\lambda}\|\leq\sqrt{m}M_{\boldsymbol{\lambda}}$, $\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}},\boldsymbol{\lambda}^{(t)})$ is $(p-L)$-strongly convex and $K$-smooth in ${\mathbf{x}}$ over $\mathcal{X}$ for any $t\geq 0$, where
$K:=$ $\leavevmode\nobreak\ L+L\sqrt{m}M_{\boldsymbol{\lambda}}+p.$ (29)
Hence, by [35, Corollary 1], to ensure (19), it suffices to ensure
$\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})\leq\frac{\epsilon_{t}^{2}}{4K}.$ (30)
Suppose that the algorithm in [35, Section 5.1.1] with the initial point set to ${\mathbf{x}}^{(t)}$ is applied to solve $\min_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$ for $k_{t}$ steps to produce ${\mathbf{x}}^{(t+1)}$. According to (30), Assumption 3.1A and [35, Theorem 6], (19) can be ensured with complexity
$k_{t}\sim\frac{1}{2}\left[\ln\left(1+\sqrt{\frac{p-L}{2(K-(p-L))}}\right)\right]^{-1}\ln\left(\frac{(K-(p-L))D_{\mathcal{X}}^{2}}{\epsilon_{t}^{2}/(4K)}\right)=O\left(\sqrt{\frac{K}{p-L}}\ln(\epsilon_{t}^{-1})\right).$ (31)
Similar complexity is needed by other accelerated FOMs, e.g. the one in [36, (2.2.22)].
@@ 5 Convergence Analysis for the iMELa Method
For simplicity of notation, let $\sigma=\frac{p-L}{p}.$ The following lemma from [50, Lemma 3.5] characterizes the Lipschitz continuity of ${\mathbf{x}}({\mathbf{z}})$ in (14) and ${\mathbf{x}}(\boldsymbol{\lambda},{\mathbf{z}})$ in (17) under Assumption 3.1.
@@
$\|{\mathbf{z}}-{\mathbf{z}}^{\prime}\|\geq$ $\leavevmode\nobreak\ \sigma\|{\mathbf{x}}({\mathbf{z}})-{\mathbf{x}}({\mathbf{z}}^{\prime})\|,$ (32)
$\|{\mathbf{z}}-{\mathbf{z}}^{\prime}\|\geq$ $\leavevmode\nobreak\ \sigma\|{\mathbf{x}}(\boldsymbol{\lambda},{\mathbf{z}})-{\mathbf{x}}(\boldsymbol{\lambda},{\mathbf{z}}^{\prime})\|.$
Similar to the analysis in [50, 51, 52], consider the potential function
$\phi^{t}=\phi({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)}):=\mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)})-2d(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)})+2v({\mathbf{z}}^{(t)}).$ (33)
By Assumption 3.1B, (13), (17) and (18), it is easy to see that
$\phi^{t}=$ $\leavevmode\nobreak\ (\mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)})-d(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)}))+(v({\mathbf{z}}^{(t)})-d(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)}))+v({\mathbf{z}}^{(t)})$
$\geq$ $\leavevmode\nobreak\ v({\mathbf{z}}^{(t)})\geq\underline{f}.$ (34)
Then, we present three descent lemmas similar to the ones in [50, 51, 52].
Lemma 5.2.
$\leavevmode\nobreak\ \mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)})-\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t+1)},\boldsymbol{\lambda}^{(t+1)})$
$\geq$ $\leavevmode\nobreak\ \frac{p-L}{2}\|{\mathbf{x}}^{(t)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}-\langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle$
Proof.
According to the definition of $\mathcal{L}_{p}$ in (16), we have
By the $(p-L)$-strong convexity of $\mathcal{L}_{p}$ in ${\mathbf{x}}$, we have
Moreover, we have from Step 6 of Algorithm 1 that
$\leavevmode\nobreak\ \mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t+1)},\boldsymbol{\lambda}^{(t+1)})$
$=$ $\leavevmode\nobreak\ \frac{p}{2}(\|{\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}-\|{\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t+1)}\|^{2})$
$=$ $\leavevmode\nobreak\ \frac{p}{2}\langle{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)},({\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)})+({\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t+1)})\rangle$
$=$ $\leavevmode\nobreak\ \frac{p}{2}(2/\theta_{t}-1)\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}$ (38)
for $\theta_{t}\leq 1$. Combining (22), (36), (37) and (38) yields (35). ∎
Lemma 5.3.
$d(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})-d(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)})\geq$ $\leavevmode\nobreak\ \langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\rangle$
Proof.
$d(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t)})-d(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)})=$ $\leavevmode\nobreak\ \mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}({\mathbf{x}}(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)})$
$\geq$ $\leavevmode\nobreak\ \mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)})$
$=$ $\leavevmode\nobreak\ \langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\rangle.$
Next, using the same technique, we have
$\leavevmode\nobreak\ d(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})-d(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t)})$
$=$ $\leavevmode\nobreak\ \mathcal{L}_{p}({\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)}),{\mathbf{z}}^{(t+1)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
$\geq$ $\leavevmode\nobreak\ \mathcal{L}_{p}({\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)}),{\mathbf{z}}^{(t+1)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}({\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
$=$ $\leavevmode\nobreak\ \frac{p}{2}(\|{\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})-{\mathbf{z}}^{(t+1)}\|^{2}-\|{\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})-{\mathbf{z}}^{(t)}\|^{2})$
$=$ $\leavevmode\nobreak\ \frac{p}{2}\langle{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)},{\mathbf{z}}^{(t+1)}+{\mathbf{z}}^{(t)}-2{\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})\rangle.$
Combining the above inequalities yields (39). ∎
Lemma 5.4.
Proof.
$\|\nabla v({\mathbf{z}})-\nabla v({\mathbf{z}}^{\prime})\|=$ $\leavevmode\nobreak\ \|p({\mathbf{z}}-{\mathbf{x}}({\mathbf{z}}))-p({\mathbf{z}}^{\prime}-{\mathbf{x}}({\mathbf{z}}^{\prime}))\|$
$\leq$ $\leavevmode\nobreak\ p(\|{\mathbf{z}}-{\mathbf{z}}^{\prime}\|+\|{\mathbf{x}}({\mathbf{z}})-{\mathbf{x}}({\mathbf{z}}^{\prime})\|)\leq p(1/\sigma+1)\|{\mathbf{z}}-{\mathbf{z}}^{\prime}\|,$
Following the analysis in [50, 51, 52], we then use the above three descent lemmas to show the decrease of the potential function $\phi^{t}$ in (33).

### v2

@@ 3 Preliminaries
$\hskip-3.61371pt\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}},\boldsymbol{\lambda}):=f({\mathbf{x}})+{\textstyle\sum_{i=1}^{m}\lambda_{i}g_{i}({\mathbf{x}})}+\frac{p}{2}\|{\mathbf{x}}-{\mathbf{z}}\|^{2}.$ (19)
Similar to the notations in Zhang and Luo [53, 54], Zhang et al. [55], we define
$d(\boldsymbol{\lambda},{\mathbf{z}}):=$ $\penalty 10000\ \min_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}},\boldsymbol{\lambda}),$ (20)
${\mathbf{x}}(\boldsymbol{\lambda},{\mathbf{z}}):=$ $\penalty 10000\ \mathop{\mathrm{arg\,min}{}}_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}},\boldsymbol{\lambda}).$ (21)
By the strong duality (see, e.g. Rockafellar [41, Sec. 28–30]), it holds that
@@ 4 Inexact Moreau Envelope Lagrangian Method
We present the inexact Moreau envelope Lagrangian (iMELa) method, shown in Algorithm 1, for finding an $\epsilon$-KKT point of (1). At iteration $t$, the method first updates the dual variable $\boldsymbol{\lambda}^{(t+1)}$ via a projected gradient ascent step on function $\mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda})$ at $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{(t)}$ with step-size $\tau_{t}$. Then, given the strongly convex subproblem $\min_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$, we denote its optimal solution by
$\tilde{{\mathbf{x}}}^{(t+1)}:={\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t)})$
in the rest of the paper for simplicity of notation. We approximately solve this subproblem by computing an $\epsilon_{t}$-stationary point ${\mathbf{x}}^{(t+1)}\in\mathcal{X}$ with $\epsilon_{t}\in(0,1)$, satisfying
@@
Such a point can be computed using the accelerated projected gradient (APG) method by Nesterov [37, Eq. (2.2.63)] with complexity provided shortly. Moreover, condition (23) can be numerically verified and thus used as a stopping criterion for the APG method, provided that $\mathcal{N}_{\mathcal{X}}({\mathbf{x}})$ has a simple structure. Finally, the proximal center $\mathbf{z}$} is updated via a convex combination of ${\mathbf{z}}^{(t)}$ and ${\mathbf{x}}^{(t+1)}$ with a weight $\theta_{t}\in[0,1]$, ensuring ${\mathbf{z}}^{(t)}\in\mathcal{X}$ for all $t\geq 0$.
@@
3:for iteration $t=0,1,\dots,T-1$ do
4:  $\lambda_{i}^{(t+1)}=[\lambda_{i}^{(t)}+\tau_{t}\cdot g_{i}({\mathbf{x}}^{(t)})]_{+},\;\forall\,i\in[m]$.
5:  ${\mathbf{x}}^{(t+1)}\approx\tilde{{\mathbf{x}}}^{(t+1)}:=\mathop{\mathrm{arg\,min}{}}_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$ that satisfies (23).
6:  ${\mathbf{z}}^{(t+1)}={\mathbf{z}}^{(t)}+\theta_{t}\cdot({\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)})$.
@@
At iteration $t$ of the iMELa method, the complexity for the APG method by Nesterov [37, Eq. (2.2.63)] to ensure (23) increases with the smoothness parameter of $\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$, which is $L+L\|\boldsymbol{\lambda}^{(t+1)}\|_{1}+p$. Fortunately, the lemma below shows that a uniform upper bound on $\{\|\boldsymbol{\lambda}^{(t)}\|\}_{t\geq 0}$ is available, whose proof is given in Section B.2.
@@
The sequence $\{\boldsymbol{\lambda}^{(t)}\}_{t\geq 0}$ generated by Algorithm 1 satisfies, for any $t\geq 0$,
$\|\boldsymbol{\lambda}^{(t)}\|\leq M_{\boldsymbol{\lambda}}:=\max\left\{2\overline{\tau}B_{g},\frac{2\overline{\tau}(C_{\boldsymbol{\lambda}}+\overline{\tau}B_{g}^{2})}{2\underline{\tau}\min_{i\in[m]}[-g_{i}({\mathbf{x}}_{\textup{feas}})]}\right\},$ (24)
@@
Thanks to Lemma 4.1 and the fact that $\|\boldsymbol{\lambda}^{(t+1)}\|_{1}\leq\sqrt{m}\|\boldsymbol{\lambda}^{(t+1)}\|\leq\sqrt{m}M_{\boldsymbol{\lambda}}$, function $\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$ is (pL)(p-L)-strongly convex and $K$-smooth in $\mathbf{x}$} over $\mathcal{X}$ for any $t\geq 0$, where
$K:=$ $\penalty 10000\ L+L\sqrt{m}M_{\boldsymbol{\lambda}}+p\geq L+L\|\boldsymbol{\lambda}^{(t+1)}\|_{1}+p.$ (25)
Given this result, the complexity for the APG method by Nesterov [37, Eq. (2.2.63)] to ensure (23) is $O(\ln(\epsilon_{t}^{-1}))$. This is stated in the following lemma whose proof is given in Section B.3.
Lemma 4.2.
Suppose the APG method by Nesterov [37, Eq. (2.2.63)] is applied to the subproblem $\min_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$ for $k_{t}$ iterations with any initial solution in $\mathcal{X}$ and returns a solution ${\mathbf{u}}^{(k_{t})}$. Let
${\mathbf{x}}^{(t+1)}=\textup{proj}_{\mathcal{X}}\big({\mathbf{u}}^{(k_{t})}-K^{-1}\cdot\nabla\mathcal{L}_{p}({\mathbf{u}}^{(k_{t})},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})\big).$
Then ${\mathbf{x}}^{(t+1)}$ satisfies (23) with
$k_{t}=O\left(\sqrt{\frac{K}{p-L}}\ln(\epsilon_{t}^{-1})\right).$
In other words, (23) is ensured with complexity $O(\ln(\epsilon_{t}^{-1}))$ for any $t\geq 0$.
The procedure described in Lemma 4.2 is formally stated in Algorithm 2 in the appendix and is implemented in our numerical experiments.
@@ 5 Convergence Analysis for the iMELa Method
For simplicity of notation, let
$\sigma=(p-L)/p.$
The following lemma from Zhang and Luo [53, Lemma 3.5] characterizes the Lipschitz continuity of ${\mathbf{x}}({\mathbf{z}})$ in (17) and ${\mathbf{x}}(\boldsymbol{\lambda},{\mathbf{z}})$ in (20) under Assumption 3.1.
@@
$\|{\mathbf{z}}-{\mathbf{z}}^{\prime}\|\geq$ $\penalty 10000\ \sigma\|{\mathbf{x}}({\mathbf{z}})-{\mathbf{x}}({\mathbf{z}}^{\prime})\|,$ (26)
$\|{\mathbf{z}}-{\mathbf{z}}^{\prime}\|\geq$ $\penalty 10000\ \sigma\|{\mathbf{x}}(\boldsymbol{\lambda},{\mathbf{z}})-{\mathbf{x}}(\boldsymbol{\lambda},{\mathbf{z}}^{\prime})\|.$ (27)
Similar to the analysis in Zhang and Luo [53, 54], Zhang et al. [55], consider the potential function
$\hskip-7.22743pt\phi^{t}:=\mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)})-2d(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)})+2v({\mathbf{z}}^{(t)}).$ (28)
By Assumption 3.1B, (16), (20) and (22), we have
$\phi^{t}=$ $\penalty 10000\ (\mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)})-d(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)}))$
$+(v({\mathbf{z}}^{(t)})-d(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)}))+v({\mathbf{z}}^{(t)})$
$\geq$ $\penalty 10000\ v({\mathbf{z}}^{(t)})\geq\underline{f}.$ (29)
Following the analysis in Zhang and Luo [53, 54], Zhang et al. [55], we characterize the change of $\phi^{t}$ after each iteration as follows. The proof is given in Section B.4.